OpenStudy (anonymous):

A show salesman marked up a pair of sandals in the spring by 75% and advertised them for $42.00. In the fall, he put them on sale at a 30% discount. What was the original price of the sandals, before the markup? How much were they selling for after the fall discount?

OpenStudy (anonymous):

Ok from what you know, you need to find out what 1/4 of 42 is. And you must find out what 30% of 42 is.

OpenStudy (alienium):

a = 75/100 of x = $42 price after fall = a + 30% of a

OpenStudy (anonymous):

Can yall show me some work? e.e

OpenStudy (anonymous):

I cannot tell you the answer but, use your brain, what is 42 divided by 4?

OpenStudy (anonymous):

10.5

OpenStudy (anonymous):

ok. so translated into money how much is that?

OpenStudy (anonymous):

$52.50?

OpenStudy (anonymous):

No. dont add 42 to it.

OpenStudy (anonymous):

anyways i need work.. i dont need to be asken the questions.....

OpenStudy (anonymous):

Ok. ill answer the first half. $10.50 is how much it was.

OpenStudy (anonymous):

i dont need answers i need the work e.e

OpenStudy (anonymous):

Shown, written down e.e..

OpenStudy (anonymous):

Ok. $42.00 * 0.25= $10.50

OpenStudy (anonymous):

Better?

OpenStudy (anonymous):

whats *?

OpenStudy (anonymous):

I need the whole questions -.-

OpenStudy (anonymous):

not pieces..

OpenStudy (anonymous):

theres 2 questions in it whats the 1st and the 2nd in work..

OpenStudy (anonymous):

I will only Tell you the first half, from the way i told you to figure it out, how would you do $42.00/30%?

OpenStudy (anonymous):

And once you do that, enter both processes toghether as your answer.

OpenStudy (anonymous):

why r u asking me questions..

OpenStudy (anonymous):

when im the one posting one...

OpenStudy (anonymous):

Its against the rules to tell you the answer.

OpenStudy (anonymous):

yet everybody tells answers on here..

OpenStudy (anonymous):

Directly? or do they tell them how to get the answers first.

OpenStudy (anonymous):

directly.

OpenStudy (anonymous):

Hmm... then they just arent being reported.

OpenStudy (anonymous):

nobody get reported for answers...

OpenStudy (anonymous):

even moderators give me answers..

OpenStudy (anonymous):

But i will give you the answer.

OpenStudy (anonymous):

lol just show the work? with answers? :L

OpenStudy (anonymous):

$42.00/25%= $10.50 , $42.00* 30%= $29.40 I hope im right

OpenStudy (anonymous):

40*30% is 12.6

OpenStudy (anonymous):

42*

OpenStudy (anonymous):

Oops, 42.00*30%= 12.6, then 42-12.6=29.40

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

42.00*30%= 12.6, then 42-12.6=29.40 makes no sense..

OpenStudy (anonymous):

the last part..

OpenStudy (anonymous):

ok, he deducted the price by 30 percent, which is 12.6 so you have to do 42-12.6

OpenStudy (anonymous):

That good?

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

nah..

OpenStudy (anonymous):

show it step by step :L

OpenStudy (anonymous):

ugh gotta go to bed... can you do it alone

OpenStudy (littlebird):

This doesn't seem right. If you increase something by 75%, doesn't that mean you need you are adding the original price and seventy five percent of the original price to get 42 dollars. That means x+.75x=42

OpenStudy (anonymous):

can you put it in all numbrs?

OpenStudy (anonymous):

instead of writing.. this is math e,e

OpenStudy (littlebird):

For the second problem, I am unsure if the number being marked down by 30% is thè original or the price after it was marked up. Either way, multiply whatever price by .30 and then subtract that value from what you started with.for example, 30% less of $42 would be $42-(.3x42)